# Large Language Model Unlearning via Embedding-Corrupted Prompts

### v1

@@ Abstract
Large language models (LLMs) have advanced to encompass extensive knowledge across diverse domains. Yet controlling what a large language model should not know is important for ensuring alignment and thus safe use. However, accurately and efficiently unlearning knowledge from an LLM remains challenging due to the potential collateral damage caused by the fuzzy boundary between retention and forgetting, and the large computational requirements for optimization across state-of-the-art models with hundreds of billions of parameters. In this work, we present Embedding-COrrupted (ECO) Prompts, a lightweight unlearning framework for large language models to address both the challenges of knowledge entanglement and unlearning efficiency. Instead of relying on the LLM itself to unlearn, we enforce an unlearned state during inference by employing a prompt classifier to identify and safeguard prompts to forget. We learn corruptions added to prompt embeddings via zeroth order optimization toward the unlearning objective offline and corrupt prompts flagged by the classifier during inference. We find that these embedding-corrupted prompts not only lead to desirable outputs that satisfy the unlearning objective but also closely approximate the output from a model that has never been trained on the data intended for forgetting. Through extensive experiments on unlearning, we demonstrate the superiority of our method in achieving promising unlearning at nearly zero side effects in general domains and domains closely related to the unlearned ones. Additionally, we highlight the scalability of our method to 100 LLMs, ranging from 0.5B to 236B parameters, incurring no additional cost as the number of parameters increases.
Figure 1: Using embedding-corrupted prompts to maintain an unlearned state on the LLM subject to unlearning We first employ a classifier to identify whether the incoming prompt falls within the scope of the unlearning target. We construct embedding-corrupted prompts by selectively corrupting dimensions within the tokens’ embeddings. The corruption parameter is learned offline via zeroth order optimization. An unlearned state is imposed during inference and does not require any updates to the original model’s weights.
@@ 1 Introduction
The use of large language models (LLMs), trained on extensive text corpora [2, 42, 6, 124, 61, 8, 137], has increasingly become standard in daily life since the arrival of ChatGPT [98]. Despite the benefits LLMs offer, they pose potential risks across a range of domains, such as copyright infringement [62, 46, 69], dissemination of hazardous knowledge [70, 50, 37, 109], and privacy violations [119, 90, 95]. Adherence to the General Data Protection Regulation (GDPR) [35], which requires the removal of users’ data post-training. Machine unlearning has emerged as a new paradigm [18, 96] and has been widely studied for classification models and tasks in recent years [123, 76, 66, 36]. However, unlearning in the context of LLMs remains largely underexplored, presenting unique challenges and risks that extend beyond privacy concerns due to the infeasibility of retraining from scratch [18, 19], the ease for anyone to access powerful models, and the substantial capabilities of these models across various tasks [77, 85].
@@
Various machine unlearning methods have been proposed specifically for LLMs to address the above challenges. A major line of approaches focuses on parameter fine-tuning [57] based on a modified loss, usually by unlearning on the forget data and learning from the retained data to preserve utility [130, 136, 22, 135, 70, 142, 59], which require only a small number of weight updates compared to retraining from scratch. Other approaches include model editing techniques [55, 132, 13, 141, 53, 97, 80], unlearning via in-context examples in the prompt [101, 94], and guarding the prompts themselves [122]. Although effective, some approaches have been shown to impair a model’s general capabilities [47, 85]. This is due to knowledge entanglement [87, 85] caused by either the fuzzy boundary between retention and forgetting objectives (e.g., forgetting a single person without affecting other related ones). Additionally, most prior work targets unlearning at the million- to billion-parameter scale through gradient-based optimization [136, 53, 34, 135, 87, 142, 70, 59], making the cost of unlearning scale with the model size and can be expensive even with parameter-efficient modules.
@@
In this work, we explore if an “unlearned state” can be imposed on an intact LLM and focus on tackling the challenges of knowledge entanglement and unlearning inefficiency in LLMs. We hypothesize that unlearning can be implemented as a state by decomposing the unlearning problem into two more tractable subproblems: 1) unlearning target identification, which explicitly identifies if the prompt contains content within the unlearning target, and 2) forgetting, which ensures that the generated responses no longer reflect any prior knowledge related to the unlearning target. We present Embedding-COrrupted (ECO [^1] Pronounced as “echo.”) Prompts, a lightweight two-step framework to tackle both problems above:
@@
    To achieve forgetting, we approximate an unlearned state by passing the query identified by the prompt classifier to the LLM, but in a corrupted form. We leverage corruptions learned efficiently via zeroth order optimization [117, 118] and apply them to the prompt’s embedding space during inference.
@@
  • We demonstrate that, instead of relying on unlearning objective optimization, carefully corrupted prompts leads to behavior that resembles that of a model which has never seen the data intended to be forgotten, across multiple tasks and metrics.
@@ 2.1 Threat Model
In our threat model, we consider threats in three categories: entity leaking, hazardous knowledge, and copyrighted content extraction. We consider a gray box setting similar to that of [70] and [122], where users interact with an LLM through a chat interface or structured API access [115], such as the GPT-4 API [2]. Under this setting, all users can send prompts to the LLM and receive the corresponding completions or per-token logits of the output tokens. We also assume that adversaries within the user group generate prompts in-distribution and attempt to jailbreak either the guarding mechanism or the LLM itself. Our threats and goals below are as follows.
@@ 2.2 Problem Setup
We assume a learning algorithm $A$ [^2] This algorithm $A$ may not be deterministic, and is assumed to be randomized., the training set $D_{tr}$ and the forget set $D_{f}$. For each dataset $D$, we have $D=\{\mathbf{z}_{i}\}_{i=1}^{N}$, where each $\mathbf{z}_{i}=\{\mathbf{x}_{i},\mathbf{y}_{i}\}$. In the traditional setting of machine unlearning, [18, 96], a retained model $\bm{\theta}_{r}$ that has never seen the forget dataset is obtained via the learning algorithm but excluding the forget set, $\bm{\theta}_{r}=A(D_{tr}\setminus D_{f})$, where $D_{r}=D_{tr}\setminus D_{f}$ is known as the retain set. We use $\bm{\theta}_{o}$ to denote the original model [^3] Throughout the paper, we also call $\bm{\theta}_{o}$ “the model subject to unlearning.” obtained from the learning algorithm $A$ and $\bm{\theta}_{r}$ to represent a retained model retrained from scratch via an unlearning algorithm $U$, which we define below, via training on on $D_{tr}$ and $D_{r}$, respectively.
Based on our threat model in Section 2.1, which does not allow users to access model weights, instead of achieving unlearning in the weight space [96], we focus on weak unlearning [11] in the output space. Specifically, we aim at similarity between models $h(\mathbf{x};\bm{\theta}_{r})$ and $h(\mathbf{x};\bm{\theta}_{u})$ for all $\mathbf{x}$, where $h:\mathcal{X}\times\Theta\rightarrow\mathcal{Y}$ maps from the input space $\mathcal{X}$ and weight space $\Theta$ to the output space $\mathcal{Y}$.
A relaxed objective of unlearning Because we are in the LLM setting, we use a relaxed definition of unlearning that does not require differential private requirement (i.e., ($\epsilon,\delta$)-close), similar to [112]. More specifically, we follow prior work [43, 22, 66, 58, 135, 51] and evaluate if the retained model and the unlearned model’s metric values over a set of metrics =$\mathcal{M}=\{m_{1},m_{2},...,m_{K}\}$ are similar, on both $D_{r}$ and $D_{f}$. To maintain the general utility of the LLM after unlearning, we would also like the model to perform well on an o.o.d. general domain distribution $\mathcal{D}_{g}$, which is unknown during unlearning. Therefore, our goal of unlearning is
@@
for all $m_{i}\in\mathcal{M}$, where $\mathcal{M}$ is a set of non-negative metrics. We want this to hold separately for each case $\mathbf{x}\sim p_{\mathcal{D}_{f}}(\mathbf{x})$, $\mathbf{x}\sim p_{\mathcal{D}_{r}}(\mathbf{x})$, and $\mathbf{x}\sim p_{\mathcal{D}_{g}}(\mathbf{x})$. During evaluation, we assess if the two models have empirically similar performance over the metrics set $\mathcal{M}$.
@@ 3.1 Method Overview
Enforcing retaining and forgetting via a classifier We first train a prompt classifier to explicitly identify if the prompt falls withing the scope of unlearning. For any incoming prompt, $\mathbf{x}$, the prompt classifier $C$ takes in $\mathbf{x}$ and returns $p_{C}(f\mid\mathbf{x})$ $=1-p_{C}(r\mid\mathbf{x})$, the probability of the prompt being in the scope of forgetting. Similar to any classifier prediction, if $p_{C}(f\mid\mathbf{x})>p_{C}(r\mid\mathbf{x})$, we consider $\mathbf{x}$ as containing the unlearning concept that our LLM is supposed to forget. Formally, given a positive prediction, $p_{C}(f\mid\mathbf{x})>p_{C}(r\mid\mathbf{x})$, we replace the original input $\mathbf{x}$ by a $\tilde{\mathbf{x}}$. Otherwise, the original $\mathbf{x}$ is passed to the LLM.
@@
Let $\tilde{h}:\mathcal{E}\times\Theta\rightarrow\mathcal{Y}$ be the function $h$ but taking the input embeddings instead of input tokens (i.e. $h$ with the input embedding layer detached), our objective is to pick a good $\sigma^{*}$ such that the following modified unlearning objective is satisfied:
@@ 3.2 Decision Threshold Calibration and Conformal Prediction
Due to the potential fuzzy boundary between retaining and forgetting, one needs to pick a threshold that is better than $p(f\mid\mathbf{x})>p(r\mid\mathbf{x})$, which does not take into account the classifier’s confidence. Depending on the application and the empirical performance of the classifier predictions, we incorporate two types of thresholding techniques.
@@
Conformal prediction We also consider conformal prediction (CP) [127], which finds a calibrated threshold given a target error rate $\alpha$, as a second way for threshold calibration. In essence, a conformal prediction uses a small user-specified error rate, $\alpha$, and unlikelihood scores (e.g., $1-p_{C}(y\mid\mathbf{x})$) on a calibration set to derive a threshold, and labels with unlikelihood scores lower than the threshold are included in the final prediction set.
We adapt the split conformal prediction setup [127], which uses a separate calibration set, $D_{\textrm{cal}}=\{\mathbf{x}_{i},y_{i}\}_{i=1}^{n}$ ($y\in\{\textrm{r},f\}$), to determine a conformity threshold, and a non-conformity score, $S:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$, to measure how unlikely a sample $(x,y)$ is to the classifier $C$. We follow convention choice and use $s_{i}=S(\mathbf{x}_{i},y_{i})=1-p_{C}(y_{i}\mid\mathbf{x}_{i})$ as the non-conformity score. Given the calibration set size $n$, and a small user-specified error rate $\alpha$, we determine a quantile $\hat{q}$ using the $\lceil(n+1)\cdot(1-\alpha)\rceil/n$ empirical quantile in the non-conformity scores from $D_{\textrm{cal}}$. The final prediction set on a new test sample $\mathbf{x}_{\textrm{test}}$ is formed by including all labels with a non-conformity score below $\hat{q}$ as
@@ 3.3 Embedding-Corrupted Prompts
Given an accurate classifier, one can already mitigate the risk defined in our threat model by providing a template response. However, doing so violates the weak unlearning objective for $\mathbf{x}\sim p_{\mathcal{D}_{f}}(\mathbf{x})$ in Equation 1, because a retained model (that is, a model not trained on the forget data) is highly unlikely to give the template responses to prompts in the forget data. To actually achieve unlearning given the prompt classifier obtained in Section 3.2, we introduce a simple method that learns to corrupt user’s prompts in the embedding space efficiently via zeroth order optimization [117, 118] toward the unlearning objective. One may also set $\sigma$ manually without optimization, at the cost of being further away from desired retrained model (see below).
Optimization objective A natural choice to make the unlearned model to behave like a retained model is to minimize a distance function which quantifies the gap between the two models for all $m\in\mathcal{M}$. As the retained model is not available (otherwise we do not need to unlearn) in practice, we use a surrogate metric value $\hat{v}_{r}$ if available to approximate how the retained model would behave over $\mathcal{M}$. Based on our relaxed unlearning objective in Equation 4, we define a general distance measure below:
$d(\tilde{\mathbf{e}},\bm{\theta}_{o},\hat{v}_{r},\mathcal{M})$ $=\frac{1}{|\mathcal{M}|}\sum_{i}\Big{|}\underbrace{m_{i}(\tilde{h}(\tilde{\mathbf{e}};\bm{\theta}_{o}))}_{\text{unlearned metric value}}-\underbrace{\hat{v}_{r}}_{\text{surrogate retain metric value}}\Big{|}$ (8)
@@
Note: If the metric value $\hat{v}_{r}$ is not possible to obtain, one may tune $\sigma$ directly and inspect if the model output on the forget set aligns with the unlearning criteria. For classification-style tasks, the target $\hat{v}_{r}$ may be known to be random guessing.
Corruption learning via zeroth order optimization We now formulate the zeroth order gradient approximation via finite differences [117, 118]. Given a pre-defined perturbation size $\mu$ applied to the current corruption parameter $\sigma_{k}$, we treat the distance function $d(\cdot)$ as a black-box and query it for the final metric gap during optimization. Because we only learn the strength of the corruption function with s scalar-valued $\sigma$, we use a deterministic perturbation to $\sigma_{k}$. For a single sample, given an initial guess $\sigma_{0}$, a step size $\eta$, a smoothing parameter $\mu$ (also known as perturbation size), the minimization of Equation 9 uses the following update rules:
@@
Choice of corruption function Prior work [38] suggests that only a small number of dimensions for each embedding vector suffices to steer the output, so we only corrupt the first dimension of each token’s embedding. We also experimented with other corruption functions (e.g., standard Gaussian noise or zeroing-out top entries), but find our method is insensitive to the corruption function and all corruption functions give similar end results. We conducted ablation studies on various corruption functions in Section D.2.

### v2

@@ Abstract
Large language models (LLMs) have advanced to encompass extensive knowledge across diverse domains. Yet controlling what a large language model should not know is important for ensuring alignment and thus safe use. However, accurately and efficiently unlearning knowledge from an LLM remains challenging due to the potential collateral damage caused by the fuzzy boundary between retention and forgetting, and the large computational requirements for optimization across state-of-the-art models with hundreds of billions of parameters. In this work, we present Embedding-COrrupted (ECO) Prompts, a lightweight unlearning framework for large language models to address both the challenges of knowledge entanglement and unlearning efficiency. Instead of relying on the LLM itself to unlearn, we enforce an unlearned state during inference by employing a prompt classifier to identify and safeguard prompts to forget. We learn corruptions added to prompt embeddings via zeroth order optimization toward the unlearning objective offline and corrupt prompts flagged by the classifier during inference. We find that these embedding-corrupted prompts not only lead to desirable outputs that satisfy the unlearning objective but also closely approximate the output from a model that has never been trained on the data intended for forgetting. Through extensive experiments on unlearning, we demonstrate the superiority of our method in achieving promising unlearning at nearly zero side effects in general domains and domains closely related to the unlearned ones. Additionally, we highlight the scalability of our method to 100 LLMs, ranging from 0.5B to 236B parameters, incurring no additional cost as the number of parameters increases. We have made our code publicly available at https://github.com/chrisliu298/llm-unlearn-eco.
Figure 1: Using embedding-corrupted prompts to maintain an unlearned state on the LLM subject to unlearning. We first employ a classifier to identify whether the incoming prompt falls within the scope of the unlearning target. We construct embedding-corrupted prompts by selectively corrupting dimensions within the tokens’ embeddings. The corruption parameter is learned offline via zeroth order optimization. An unlearned state is imposed during inference and does not require any updates to the original model’s weights.
@@ 1 Introduction
The use of large language models (LLMs), trained on extensive text corpora [2, 45, 6, 131, 64, 8, 144], has increasingly become standard in daily life since the arrival of ChatGPT [103]. Despite the benefits LLMs offer, they pose potential risks across a range of domains, such as copyright infringement [65, 49, 72], dissemination of hazardous knowledge [73, 53, 39, 115], and privacy violations [126, 95, 100]. Adherence to the General Data Protection Regulation (GDPR) [37], which requires the removal of users’ data post-training, is essential. Machine unlearning has emerged as a new paradigm [19, 101] and has been widely studied for classification models and tasks in recent years [130, 79, 69, 38]. However, unlearning in the context of LLMs remains largely underexplored, presenting unique challenges and risks that extend beyond privacy concerns due to the infeasibility of retraining from scratch [19, 20], the ease with which anyone can access powerful models, and the substantial capabilities of these models across various tasks [80, 88].
@@
Various machine unlearning methods have been proposed specifically for LLMs to address the above challenges. A major line of approaches focuses on parameter fine-tuning [60] based on a modified loss, usually by unlearning on the forget data and learning from the retained data to preserve utility [137, 143, 23, 142, 73, 149, 62], which requires only a small number of weight updates compared to retraining from scratch. Other approaches include model editing techniques [58, 139, 13, 148, 56, 102, 83], unlearning via in-context examples in the prompt [106, 99], and guarding the prompts themselves [129]. Although effective, some approaches have been shown to impair a model’s general capabilities [50, 88]. This is due to knowledge entanglement caused by the fuzzy boundary between retention and forgetting objectives (e.g., forgetting a single person without affecting other related ones) [93, 90, 88, 109]. Additionally, most prior work targets unlearning at the million- to billion-parameter scale through gradient-based optimization [143, 56, 36, 142, 90, 149, 73, 62], making the cost of unlearning scale with the model size and potentially expensive even with parameter-efficient modules. This cost could rise significantly for advanced proprietary models with hundreds of billions of parameters (e.g., GPT-4 [2], Gemini [45], Claude [6], and other model-as-a-service (MaaS) [42] providers), which makes gradient-based unlearning methods impractical.
@@
In this work, we explore if an “unlearned state” can be imposed on an intact LLM and focus on tackling the challenges of knowledge entanglement and unlearning inefficiency in LLMs. We hypothesize that unlearning can be implemented as a state by decomposing the unlearning problem into two more tractable subproblems: 1) unlearning target identification, which explicitly identifies if the prompt contains content within the unlearning target, and 2) forgetting, which ensures that the generated responses no longer reflect any prior knowledge related to the unlearning target. We present Embedding-COrrupted (ECO) Prompts, a lightweight two-step framework to tackle both problems above:
@@
    To achieve forgetting, we approximate an unlearned state by passing the query identified by the prompt classifier to the LLM, but in a corrupted form. We leverage corruptions learned efficiently via zeroth-order optimization [124, 125] and apply them to the prompt’s embedding space during inference.
@@
  • We demonstrate that, instead of relying on unlearning objective optimization, carefully corrupted prompts lead to behavior that resembles that of a model which has never seen the data intended to be forgotten, across multiple tasks and metrics.
@@ 2.1 Threat Model
In our threat model, we consider threats in three categories: entity leaking, hazardous knowledge, and copyrighted content extraction. We consider a gray box setting similar to that of [73] and [129], where users interact with an LLM or a model-as-a-service [42] through a chat interface or structured API access [121]. Under this setting, all users can send prompts to the LLM and receive the corresponding completions or per-token logits of the output tokens. We also assume that adversaries within the user group generate prompts in-distribution and attempt to jailbreak either the guarding mechanism or the LLM itself. Our threats and goals below are as follows.
@@
Beyond the categorization of risks presented above, we also highlight a commonly overlooked aspect in unlearning: timeliness [109, 123, 16, 101]. Timeliness measures how quickly unlearning can be completed once the relevant risks are identified. Given the volume of real-time interactions from MaaS users [42], the effectiveness of LLM unlearning may degrade progressively with each hour of delay, particularly in safety and privacy domains. Our objective is to develop a method that can be implemented with extreme efficiency, ideally operating in real-time.
@@ 2.2 Problem Setup
We assume a learning algorithm $A$ [^1] This algorithm $A$ may not be deterministic and is assumed to be randomized., the training set $D_{tr}$, and the forget set $D_{f}$. For each dataset $D$, we have $D=\{\mathbf{z}_{i}\}_{i=1}^{N}$, where each $\mathbf{z}_{i}=\{\mathbf{x}_{i},\mathbf{y}_{i}\}$. In the traditional setting of machine unlearning [19, 101], a retained model $\bm{\theta}_{r}$ that has never seen the forget dataset is obtained via the learning algorithm but excluding the forget set, $\bm{\theta}_{r}=A(D_{tr}\setminus D_{f})$, where $D_{r}=D_{tr}\setminus D_{f}$ is known as the retain set. We use $\bm{\theta}_{o}$ to denote the original model [^2] Throughout the paper, we also call $\bm{\theta}_{o}$ “the model subject to unlearning.” obtained from the learning algorithm $A$, and $\bm{\theta}_{r}$ to represent a retained model retrained from scratch via an unlearning algorithm $U$, which we define below, by training on $D_{tr}$ and $D_{r}$, respectively.
Based on our threat model in Section 2.1, which does not allow users to access model weights, instead of achieving unlearning in the weight space [101], we focus on weak unlearning [11] in the output space. Specifically, we aim for similarity between models $h(\mathbf{x};\bm{\theta}_{r})$ and $h(\mathbf{x};\bm{\theta}_{u})$ for all $\mathbf{x}$, where $h:\mathcal{X}\times\Theta\rightarrow\mathcal{Y}$ maps from the input space $\mathcal{X}$ and weight space $\Theta$ to the output space $\mathcal{Y}$.
A relaxed objective of unlearning Because we are in the LLM setting, we use a relaxed definition of unlearning that does not require differential privacy requirements (i.e., ($\epsilon,\delta$)-close), similar to [118]. More specifically, we follow prior work [46, 23, 69, 61, 142, 54] and evaluate whether the retained model and the unlearned model’s metric values over a set of metrics =$\mathcal{M}=\{m_{1},m_{2},...,m_{K}\}$ are similar on both $D_{r}$ and $D_{f}$. To maintain the general utility of the LLM after unlearning, we would also like the model to perform well on an o.o.d. general domain distribution $\mathcal{D}_{g}$, which is unknown during unlearning. Therefore, our goal of unlearning is
@@
for all $m_{i}\in\mathcal{M}$, where $\mathcal{M}$ is a set of non-negative metrics. We want this to hold separately for each case $\mathbf{x}\sim p_{\mathcal{D}_{f}}(\mathbf{x})$, $\mathbf{x}\sim p_{\mathcal{D}_{r}}(\mathbf{x})$, and $\mathbf{x}\sim p_{\mathcal{D}_{g}}(\mathbf{x})$. During evaluation, we assess whether the two models have empirically similar performance over the metrics set $\mathcal{M}$.
@@ 3.1 Method Overview
Enforcing retaining and forgetting via a classifier We first train a prompt classifier to explicitly identify if the prompt falls within the scope of unlearning. For any incoming prompt, $\mathbf{x}$, the prompt classifier $C$ takes in $\mathbf{x}$ and returns $p_{C}(f\mid\mathbf{x})=1-p_{C}(r\mid\mathbf{x})$, the probability of the prompt being in the scope of forgetting. Similar to any classifier prediction, if $p_{C}(f\mid\mathbf{x})>p_{C}(r\mid\mathbf{x})$, we consider $\mathbf{x}$ as containing the unlearning concept that our LLM is supposed to forget. Formally, given a positive prediction, $p_{C}(f\mid\mathbf{x})>p_{C}(r\mid\mathbf{x})$, we replace the original input $\mathbf{x}$ with $\tilde{\mathbf{x}}$. Otherwise, the original $\mathbf{x}$ is passed to the LLM.
@@
Let $\tilde{h}:\mathcal{E}\times\Theta\rightarrow\mathcal{Y}$ be the function $h$ but taking the input embeddings instead of input tokens (i.e., $h$ with the input embedding layer detached). Our objective is to pick a good $\sigma^{*}$ such that the following modified unlearning objective is satisfied:
@@ 3.2 Decision Threshold Calibration and Conformal Prediction
Due to the potential fuzzy boundary between retaining and forgetting, one needs to pick a threshold better than $p(f\mid\mathbf{x})>p(r\mid\mathbf{x})$, which does not take into account the classifier’s confidence. Depending on the application and the empirical performance of the classifier predictions, we incorporate two types of thresholding techniques.
@@
Conformal prediction We also consider conformal prediction (CP) [134], which finds a calibrated threshold given a target error rate $\alpha$, as a second way for threshold calibration. In essence, conformal prediction uses a small user-specified error rate, $\alpha$, and unlikelihood scores (e.g., $1-p_{C}(y\mid\mathbf{x})$) on a calibration set to derive a threshold. Labels with unlikelihood scores lower than the threshold are included in the final prediction set.
We adapt the split conformal prediction setup [134], which uses a separate calibration set, $D_{\textrm{cal}}=\{\mathbf{x}_{i},y_{i}\}_{i=1}^{n}$ ($y\in\{\textrm{r},f\}$), to determine a conformity threshold and a non-conformity score, $S:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$, to measure how unlikely a sample $(x,y)$ is to the classifier $C$. Following conventional choice, we use $s_{i}=S(\mathbf{x}_{i},y_{i})=1-p_{C}(y_{i}\mid\mathbf{x}_{i})$ as the non-conformity score. Given the calibration set size $n$ and a small user-specified error rate $\alpha$, we determine a quantile $\hat{q}$ using the $\lceil(n+1)\cdot(1-\alpha)\rceil/n$ empirical quantile in the non-conformity scores from $D_{\textrm{cal}}$. The final prediction set on a new test sample $\mathbf{x}_{\textrm{test}}$ is formed by including all labels with a non-conformity score below $\hat{q}$ as
@@ 3.3 Embedding-Corrupted Prompts
Given an accurate classifier, one can already mitigate the risk defined in our threat model by providing a template response. However, doing so violates the weak unlearning objective for $\mathbf{x}\sim p_{\mathcal{D}_{f}}(\mathbf{x})$ in Equation 1, because a retained model (that is, a model not trained on the forget data) is highly unlikely to give template responses to prompts in the forget data. To actually achieve unlearning, given the prompt classifier obtained in Section 3.2, we introduce a simple method that learns to corrupt user prompts in the embedding space efficiently via zeroth order optimization [124, 125] toward the unlearning objective. One may also set $\sigma$ manually without optimization, at the cost of being further away from the desired retained model (see below).
Optimization objective A natural choice to make the unlearned model behave like a retained model is to minimize a distance function that quantifies the gap between the two models for all $m\in\mathcal{M}$. As the retained model is not available (otherwise, unlearning would not be needed), we use a surrogate metric value $\hat{v}_{r}$ if available to approximate how the retained model would behave over $\mathcal{M}$. Based on our relaxed unlearning objective in Equation 4, we define a general distance measure below:
$d(\tilde{\mathbf{e}},\bm{\theta}_{o},\hat{v}_{r},\mathcal{M})$ $=\frac{1}{|\mathcal{M}|}\sum_{i}\Big{|}\underbrace{m_{i}(\tilde{h}(\tilde{\mathbf{e}};\bm{\theta}_{o}))}_{\text{unlearned metric value}}-\underbrace{\hat{v}_{r}}_{\text{surrogate retained metric value}}\Big{|}$ (8)
@@
Note: If the metric value $\hat{v}_{r}$ is not obtainable, one may tune $\sigma$ directly and inspect whether the model output on the forget set aligns with the unlearning criteria. For classification-style tasks, the target $\hat{v}_{r}$ may correspond to random guessing.
Corruption learning via zeroth order optimization We now formulate the zeroth order gradient approximation via finite differences [124, 125]. Given a pre-defined perturbation size $\mu$ applied to the current corruption parameter $\sigma_{k}$, we treat the distance function $d(\cdot)$ as a black-box and query it for the final metric gap during optimization. Because we only learn the strength of the corruption function with a scalar-valued $\sigma$, we use a deterministic perturbation to $\sigma_{k}$. For a single sample, given an initial guess $\sigma_{0}$, a step size $\eta$, and a smoothing parameter $\mu$ (also known as perturbation size), the minimization of Equation 9 uses the following update rules:
@@
Choice of corruption function Prior work [40] suggests that only a small number of dimensions for each embedding vector suffices to steer the output, so we only corrupt the first dimension of each token’s embedding. We also experimented with other corruption functions (e.g., standard Gaussian noise or zeroing-out top entries) but found that our method is insensitive to the choice of corruption function, with all tested functions yielding similar end results. We conducted ablation studies on various corruption functions in Section D.2.